### v1

# Josephson anomalous vortices

Abstract
We show that vortices with circulating current, related with odd-frequency triplet pairing, appear in Josephson junctions where the barrier is a weak ferromagnet with strong spin-orbit coupling. By both symmetry analysis and microscopic methods we show that there is an additional term — a rotary invariant — in the superconducting free energy which allows for magnetoelectric effects even when the previously considered Lifshitz invariant vanishes. We show that the size, shape, and position of these vortices can be controlled by manipulating Rashba spin-orbit coupling in the weak link, via gates, and we suggest that these vortices could be detected via scanning magnetometry techniques. We also show that the transverse triplet components of the superconducting correlations can form a texture.
Introduction— Vortices in superconductors have remained a vibrant area of research since their discovery in the 1950s. In the original work by Abrikosov [1], vortices appear in type II superconductors threaded by a flux quantum as a result of the superconducting phase winding by $2\pi$ around a singularity. This produces a signature circulating current with a normal core. Superconducting heterostructures allow for a rich variety of other kinds of vortices. For example, in Josephson junctions coreless, elongated vortices can arise due to kinks or solitons in the superconducting phase [2, 3]. Vortices with cores can also appear in normal metals in proximity to a superconductor [4], and also in junctions due to destructive interference of the superconducting correlations [5]. While in these cases, an external magnetic field is required to generate vortices, other forms of magnetism can also generate vortices. For example, evidence has been presented for vortices in Fe(Se, Te) due to the internal magnetization [6, 7], and on superconducting surfaces with strong spin-orbit coupling (SOC) and magnetic adatoms [8]. In the latter case, spin-triplet superconducting correlations are created by the ferromagnetic proximity effect, and modulated by spin-orbit coupling. In analogy to superfluid He-3 [9, 10, 11], it might be possible to generate also more exotic vortex matter with such triplet correlations.
Figure 1: (a) Sketch of a superconductor-ferromagnetic metal-superconductor device studied here. There is strong spin-orbit coupling in the ferromagnetic layer. (b) Circulating current density in the weak link of (a) due to the rotary invariant. Parameters: $(L_{x},L_{y},J,\alpha)=(\xi_{0},\xi_{0},\Delta_{0},2.35/\xi_{0})$.
Here we introduce the Josephson anomalous vortex (JAV). These appear as a circulating current density (Fig. missing 1(b)) in two-dimensional junctions where the Josephson weak link is ferromagnetic and has strong SOC (Fig. missing 1(a)), as long as no current is applied. JAVs only appear when the weak link, or gate-defined regions in the weak link, are close to a coherence length in length and width. The current density is at a maximum at the edges of the vortex, and vanishes at the center. Complementary to these current vortices, we also present two kinds of superconducting textures: (i) a texture in the phase of the $f_{z}$ component, and (ii) a texture formed from the amplitudes of the $f_{x}$ and $f_{y}$ components of the superconducting correlations. The latter are analogous to a magnetic texture.
In this Letter we study a previously unexplored rotary invariant in the free energy. We show that this invariant generates circulating currents, even when the standard Lifshitz invariant [12] vanishes. We show that JAVs are not fine-tuned so that they can appear over a range of spin-orbit and exchange energies. We argue that, due to the magnetic field produced by swirling currents, JAVs could be detected by scanning magnetometry techniques. We also suggest that these vortices may be manipulated by Rashba “islands” within the weak link via back gates.
Throughout this Letter we use the singlet superconducting gap $\Delta_{0}$ of the electrodes as the natural unit of energy, and the superconducting coherence length $\xi_{0}$ as the unit of length. We restore the physical units when predicting amplitudes of measurable quantities. In numerical calculations we fix the temperature to $T=0.1\Delta_{0}/k_{\rm B}$. We use the conventions that Greek subscripts $\mu,\nu$ run over $(0,1,2,3)$, while Latin subscripts run over $(1,2,3)$. Repeated indices imply summation.
Magnetoelectric effects without a Lifshitz invariant— Over the past decades, it has been shown experimentally and theoretically that SOC allows for coupling between the magnetic and electric properties of a material [12]. Such coupling is expressed in phenomena such as the superconducting and Josephson diode effects [13], helical phases [14], and the Edelstein and inverse Edelstein effects [15, 16, 17]. For weak superconductivity many magnetoelectric effects can be described by the addition of a Lifshitz invariant in the Ginzburg-Landau free energy [12]. In this Letter we investigate a rotary invariant in the free energy which also gives rise to magnetoelectric effects, even when there is no Lifshitz invariant. This term was reported in [18], but its significance was not discussed.
Consider first the superconducting free energy density
$\mathcal{F}=\,$ $a|\psi|^{2}+\frac{b}{2}|\psi|^{4}+\frac{1}{2}D_{jk}\Pi_{j}\psi(\Pi_{k}\psi)^{*}+\frac{|{\bm{B}}|^{2}}{2\mu_{0}}$
$+d_{k}[\psi^{*}\Pi_{k}\psi+\psi(\Pi_{k}\psi)^{*}]$
$+ie_{jk}\Pi_{j}\psi(\Pi_{k}\psi)^{*}.$ (1)
Here $\psi$ is the complex order parameter of the superconductor, $a$ and $b$ are the usual Ginzburg-Landau constants, $D_{jk}$ the superfluid stiffness symmetric tensor, and ${\bm{B}}$ the magnetic field. For compactness we define the momentum $\Pi_{j}=-i\hbar\hat{\nabla}_{j}=-i\hbar\partial_{j}-2eA_{j}$, with $A_{j}$ the electromagnetic vector potential. In addition to the conventional Ginzburg-Landau terms, symmetries allow several additional terms [19]: the Lifshitz invariant vector $d_{k}$, and the rotary invariant antisymmetric tensor $e_{jk}$ which is of interest here.
The presence of $d_{k}$ and $e_{jk}$ requires breaking of time-reversal symmetry. The vector $d_{k}$ also requires intrinsic breaking of inversion symmetry. The $e_{jk}$ term is a total derivative, and requires breaking of inversion symmetry by extrinsic inhomogeneity, for example system boundaries (see Appendix A for details). One mechanism for breaking those symmetries comes from the combination of SOC and exchange field ${\bm{h}}$, which we consider here. In the case of linear in momentum SOC, with $H_{\rm SO}=\mathcal{A}_{j}^{a}k_{j}\sigma^{a}/(2m)$, we can define the SU(2) fields $\mathcal{A}_{j}=\mathcal{A}_{j}^{a}\sigma^{a}$ [20, 21] and the corresponding field strength tensor $F_{ij}=-i[\mathcal{A}_{i},\mathcal{A}_{j}]$, valid for position independent SU(2) fields. Using these definitions, we find that (see Sec. S2 in [22]) generally
$d_{i}=i\kappa{\rm Tr}\left\{F_{ij}[\mathcal{A}_{j},h_{k}\sigma_{k}]\right\},\quad e_{ij}=\kappa{\rm Tr}[F_{ij}h_{k}\sigma_{k}],$ (2)
where $\kappa=-\frac{7\zeta(3)\nu_{\rm F}\eta}{16\pi^{2}k_{B}^{2}T_{c}^{2}}$, with $\nu_{\rm F}$ the density of states per spin projection at the Fermi level, $\eta=D\ell^{2}/(k_{\rm F}\ell)$, $D=v_{\rm F}\ell/2$ the diffusion constant, $v_{\rm F}$ the Fermi velocity, $k_{\rm F}$ the Fermi wavevector, $T_{c}$ is the critical temperature, and $\ell$ the mean free path. $\zeta(n)$ is the Zeta function. Thus $d_{k}$ is cubic and $e_{ij}$ quadratic in the strength of SOC. Moreover, these forms imply that $d_{k}$ is non-zero only when the exchange field and $F_{ij}$ have perpendicular components, whereas $e_{ij}$ requires the presence of their parallel components. We can hence study the implications of one without the other.
In the setup shown in Fig. missing 1(a), with an exchange field ${\bm{h}}=J\hat{{\bm{z}}}$ and Rashba SOC due to an electric field also in the $z$ direction, the SOC Hamiltonian is $H_{\rm SO}=\alpha(\hat{{\bm{z}}}\times{\bm{k}})\cdot{\bm{\sigma}}$ and the corresponding field-strength tensor is $F_{xy}=-F_{yx}=2\alpha^{2}\sigma_{z}$. Hence, $d_{k}=0$ and $e_{xy}=-e_{yx}=4\kappa\alpha^{2}J$. Integration by parts of this rotary term results in a boundary term $\propto{}n_{j}e_{jk}[\psi\partial_{k}\psi^{*}-\mathrm{c.c.}]$ that generates circulating currents perpendicular to the boundary normal $n_{j}$.
We derive the rotary invariant from microscopics in the diffusive regime, based on the methods described in Ref. 23, 24 (see Sec. S1 in [22] for details). While Eq. (missing) 1 is valid for a bulk superconducting system, for heterostructures such as the one sketched in Fig. missing 1(a) the corresponding rotary invariant of the microscopic theory is
$\mathcal{F}_{\rm rot}=i\zeta(\hat{\nabla}f_{s}\times\hat{\nabla}^{*}f_{z}^{*})\cdot\hat{{\bm{z}}}.$ (3)
Here $\zeta=\pi\nu_{\rm F}\hbar\eta\alpha^{2}$; in all numerical results we assume $\hbar\eta=0.01\Delta_{0}\xi_{0}^{4}$. As this is a proximity system we write the free energy in terms of superconducting correlations ${\bm{f}}=f_{s}\sigma_{0}+f_{i}\sigma_{i}$, and due to the ferromagnetic proximity effect we require that the correlations have both singlet $f_{s}$ and triplet ${\bm{f}}_{t}=(f_{x},f_{y},f_{z})$ components [25]. We assume that the ferromagnet is weak, so that the quasiclassical approximation is valid (i.e., the density of states is the same for both up and down spins). The rotary term $\mathcal{F}_{\rm rot}$ appears as long as there is both SOC and an exchange field, even when there is no Lifshitz invariant. It generates circulating current densities,
${\bm{J}}_{\rm rot}=2e\zeta\left(f_{s}\text{curl}\,f_{z}^{*}+f_{s}^{*}\text{curl}\,f_{z}\right),$ (4)
with $\text{curl}\,A$ = $(\partial_{y}A,-\partial_{x}A)$ and $e$ the elementary charge. Here $f_{s}$ varies across the junction due to the proximity effect and ${\bm{f}}_{t}$ is in general finite due to the presence of the exchange field. We expect the conventional Ginzburg-Landau terms to dominate over the rotary invariant, so to see the effects of this term there should be no current (i.e., phase bias $\varphi$) across the junction.
We numerically solve the full nonlinear problem, including singlet-triplet conversion and finite-size effects, using the methods described in Ref. 18. In Fig. missing 1(b) we plot the current density for a junction of $L_{x}\times L_{y}=\xi_{0}\times\xi_{0}$, at zero phase bias. This yields a representative current vortex, distinguished by the strong current density at the boundary (as opposed to the center, as in an Abrikosov vortex). For large junctions we find some swirling current close to the superconducting terminals [22], similar to what has been reported in Ref. 26 for in-plane exchange fields. As long as $\alpha^{2}\xi_{0}^{2}\gg\varphi$, the rotary invariant, and thus the circulating current, dominate over the usual Josephson current. For $\varphi\gtrsim\alpha^{2}\xi_{0}^{2}$ the rotary invariant perturbs the Josephson current (see Sec. S4 in [22]).
Figure 2: Magnetic field strength along the $\hat{z}$ direction at the center of the junction due to circulating currents for various junction sizes, over a range of exchange energies $J$ and spin-orbit coupling strengths $\alpha$. Dashed lines indicate the transitions (T) between antivortex (AV) and vortex (V) phases. Inset: representative plot of an antivortex. Parameters for inset: $(L_{x},L_{y},J,\alpha)=(\xi_{0},\xi_{0},\Delta_{0},1/(2\xi_{0}))$.
Magnetic field due to a vortex— Using the Biot-Savart law, we plot in Fig. missing 2 the magnetic field at the centre of the vortex, for four different junction sizes at zero phase bias, over a range of exchange energies $J$ and SOC strengths $\alpha$. The magnetic field scale in the plot is $B_{0}=\mu_{0}\sigma_{\rm D}\Delta_{0}t/(4e\xi_{0})=1$ mT×$\times(t/\xi_{0})[\sigma_{\rm D}/(1.6\cdot 10^{7}\text{ S/m})][\Delta_{0}/(200\text{ $\mu$eV})]$, expressed in terms of relatively typical values for the Drude conductivity $\sigma_{\rm D}=\nu_{\rm F}e^{2}D$ in disordered metals, and a typical $\Delta_{0}$ for Al. We also include the thickness $t\ll\xi_{0}$ of the junction. We do not compute this magnetic field self-consistently (i.e., by feeding back into the electromagnetic vector potential). Due to competition between the rotary invariant and other terms in the free energy, current vortices can only appear above some threshold $\alpha,J$, depending, e.g., on boundary conditions, temperature, and field strength. We show in Appendix B, for example, that from the Ginzburg-Landau free energy vortices are stable when the rotary invariant exceeds the superfluid weight. Numerically, we find a vortex- antivortex transition as a function of $\alpha$ and $J$ (we plot a representative antivortex in the inset of Fig. missing 2). The antivortex phase arises due to the boundary conditions, with the exact dependence on $\alpha$ and $J$ varying with system size. Note that in the antivortex phase there is no net current transport across the junction and no associated magnetic field.
[…]
${\bm{J}}_{\rm V}=J_{0}\left(\frac{-y}{\xi_{0}}\hat{{\bm{x}}}+\frac{x}{\xi_{0}}\hat{{\bm{y}}}\right),$ ${\bm{J}}_{\rm AV}=J_{0}\left(\frac{y}{\xi_{0}}\hat{{\bm{x}}}+\frac{x}{\xi_{0}}\hat{{\bm{y}}}\right),$ (5)
with $J_{0}=e\zeta\Delta_{0}/(\hbar\xi_{0})$. It follows then that for an antivortex the circulation is $\Gamma_{\rm AV}=0$, but for a vortex the circulation is $\Gamma_{\rm V}=2J_{0}A$, with $A$ the area of the vortex. By computing $\Gamma$ over a square patch in the center of the weak link in the vortex phase, we have that $J_{\rm max}\approx J_{0}$. Hence in the numerics $\Gamma\approx 2A$ indicates a vortex, $\Gamma\approx 0$ an antivortex, and $\Gamma\approx A$ the transitional phase. Due to finite size effects and differences between the square geometry as compared to a disc, the numerical vorticity is not quantized and we see small variations compared to the analytic result.
In polar coordinates, the approximate current density for a vortex is ${\bm{J}}_{\rm V}=J_{0}\rho/\xi_{0}\hat{{\bm{\phi}}}$, with $\rho$ the radial coordinate and $\phi$ the angular coordinate. The magnetic field at the center of the vortex ($\rho=0$) is then
${\bm{B}}_{\rm anom}=\mu_{0}J_{0}\frac{Rt}{2\xi_{0}},$ (6)
with $\mu_{0}$ the vacuum permeability, and $R$ the radius of the vortex. Substituting in $R=\xi_{0}$, we find a maximum magnetic field strength of $B_{\rm anom}=2\pi(R/\xi_{0})(\alpha\xi_{0})^{2}(\eta/\Delta_{0}\xi_{0}^{4})B_{0}$, some 1.5 times larger than what is seen in numerics.
For very small ($L_{x}\times L_{y}=0.1\xi_{0}\times 0.1\xi_{0}$) junctions the system is dominated by the antivortex phase, and so the magnetic field at the center is usually small. For large $\alpha$ the system is close to the transition and so due to finite size we find a small magnetic field. For large junctions ($L_{x}\times L_{y}=5\xi_{0}\times 5\xi_{0}$) the parameter space is dominated by the vortex phase and the transitional phase. Large counterpropagating currents around the terminals counteract the magnetic field due to the central vortex (see Sec. S3 in [22]). For $L_{x}\times L_{y}=0.5\xi_{0}\times 0.5\xi_{0}$ junctions the parameter space is evenly mixed between the antivortex, vortex, and transitional phases. At around $J=\Delta_{0}$ there is a large magnetic field due to this transition (see Sec. S3 in [22]). When the system is only a little larger ($L_{x}\times L_{y}=\xi_{0}\times\xi_{0}$), the parameter space is dominated by the vortex phase, with a small transitional phase. At around $J=\Delta_{0}$ there is a large magnetic field of around $B_{\rm anom}=0.04B_{0}$ due to the vortices, in agreement with the simple approximation above. Because $f_{z}\sim J/(J^{2}+\omega^{2})$ (with $\omega$ the Matsubara frequency) [22], we expect the magnetic field strength to peak at some finite $J$, and decay for large $J$; in the numerics, this peak occurs around $J\approx\alpha/2$.
Figure 3: The size (a), position (b), and shape (c) of a JAV could be manipulated by defining an island with strong Rashba SOC $\alpha_{\rm island}$ via gates, within a weak link with weak SOC $\alpha_{\rm back}$. Inset: sketches of the gate defining $\alpha_{\rm island}$ (yellow patch) within the weak link (blue square). Parameters: $(L_{x},L_{y},J,\alpha_{\rm island},\alpha_{\rm back})$ = $(5\xi_{0},5\xi_{0},\Delta_{0},2/\xi_{0},1/(2\xi_{0}))$.
Characterization and manipulation— Because circulating currents produce a magnetic field, a natural approach to characterizing these vortices is via nano-magnetometry. Scanning probe techniques are well-established and can operate at cryogenic temperatures with high spatial resolution. Magnetic force microscopy has already been used to study conventional Josephson vortices [27, 28], and can operate with a resolution of 10 — 50 nm [29], or even atomic resolution [30]. Scanning nitrogen-vacancy center [31] and nanoSQUID methods [32] can now both achieve spatial resolutions of around 50 nm. These magnetometry techniques are complicated by the need to subtract the background ferromagnetic signal. One possible solution is to include a back gate to control the Rashba interaction [33]. One could then effectively switch off SOC to isolate the background ferromagnetic signal. The absence of SOC can, for example, be confirmed via the absence of a Josephson diode effect.
Such gate-based techniques for controlling Rashba SOC could also be used to manipulate these vortices. Consider a large weak link which has only weak (or absent) Rashba SOC. One can define an island within the weak link with strong Rashba SOC via a small back gate. We show in Fig. missing 3 examples where such Rashba “islands” can be used to control the size, position, and shape of a JAV. Such methods could be used to realize JAVs in arbitrarily sized junctions, overcoming the necessity of defining a square geometry of a coherence length. Indeed, several different vortices could be defined within a single weak link using this technique.
Superconducting textures— All components of the superconducting correlations must respect the 180°$\degree$ rotation symmetry of the junction. This requirement, in combination with the junction geometry and spin-orbit coupling, allows for phase textures of $f_{z}$, such as shown in Fig. missing 4(a). This texture is characterized by phase gradients in both the $x$ and $y$ directions. We find similar textures for all system sizes studied, above a threshold $J,\alpha$. We also point out that the $f_{x}$ and $f_{y}$ components can be finite in superconducting heterostructures with both SOC and ferromagnetism.
Consider a ferromagnetic disk with strong SOC and a superconducting rim. Due to the rotational symmetry $f_{s}$ and $f_{z}$ can only have radial dependence. In contrast, $f_{x}$ and $f_{y}$ can have both radial and angular dependence, implying a texture could form. We solve the full problem numerically for such a disc (Fig. missing 4(b)) and also for a square geometry (Fig. missing 4(c)), such as the one studied in Figs. 1-3. Indeed, we find a striking resemblance between these $f_{x}-f_{y}$ textures and magnetic textures. It is currently unclear how to observe these vortices.
Figure 4: (a) Phase texture in a square geometry, computed numerically. At $x\approx 0$, $f_{z}=0$. Superconducting $f_{x},f_{y}$-textures computed (b) numerically on a square geometry ($L_{x}=L_{y}=\xi_{0}$), and (c) numerically on a disc geometry, with the rim superconducting and the bulk of the disc a ferromagnet with SOC (radius $R=\xi_{0}$). We plot the components at the smallest Matsubara frequency. Parameters: $(J,\alpha)=(\Delta_{0},2/\xi_{0})$.
Outlook— We describe the characterization and manipulation of Josephson anomalous vortices, which appear in Josephson junctions where the weak link is a weak ferromagnet with strong spin-orbit coupling. The exchange field and spin-orbit coupling conspire to lead to circulating currents that we associate phenomenologically to a previously unexplored rotary term in the free energy. We also find that the same ingredients lead to textures in the triplet components ${\bm{f}}_{t}$, which do not directly contribute to the current.
Josephson anomalous vortices could be found in weak links based on heterostructures such as ferromagnetic thin film/normal metal bilayers, ferromagnetic alloys deposited as thin films on metals, or two-dimensional van der Waals heterostructures, coupled to conventional superconducting electrodes. Possible thin film bilayers include YIG/InAs or EuS/InAs [34], which must be tuned so that the magnetization turns out-of-plane (see e.g., Ref.   [35] and references therein). Candidate ferromagnetic alloys include Pd/Ni [36] or Cu/Ni alloys [37]. The layers must be chosen carefully to minimize the Dzaloshinskii-Moriya interaction, which will produce a magnetic texture rather than a pure out-of-plane exchange field. Finally, we suggest that the van der Waals monolayer magnet CrI3, which has out-of-plane magnetization [38], could be combined with other van der Waals materials with strong Rashba spin-orbit coupling, such as MoS2 or WSe2 [39], to find the anomalous vortices.
Acknowledgements.
The authors acknowledge discussions with G. E. Volovik, E. Strambini, A. Dunbrack, and R. Ojajärvi. This work was supported by the European Union’s HORIZON-RIA programme (Grant Agreement No. 101135240 JOGATE), managed by the Chips Joint Undertaking, and the Research Council of Finland (Contract No. 355056 and 354735). We acknowledge grants of computer capacity from the Finnish Grid and Cloud Infrastructure (persistent identifier urn:nbn:fi:research-infras-2016072533).